### v2

# LLM3: Large Language Model-based Task and Motion Planning with Motion Failure Reasoning

Abstract
Conventional Task and Motion Planning (TAMP) approaches rely on manually crafted interfaces connecting symbolic task planning with continuous motion generation. These domain-specific and labor-intensive modules are limited in addressing emerging tasks in real-world settings. Here, we present LLM3, a novel Large Language Model (LLM)-based TAMP framework featuring a domain-independent interface. Specifically, we leverage the powerful reasoning and planning capabilities of pre-trained LLMs to propose symbolic action sequences and select continuous action parameters for motion planning. Crucially, LLM3 incorporates motion planning feedback through prompting, allowing the LLM to iteratively refine its proposals by reasoning about motion failure. Consequently, LLM3 interfaces between task planning and motion planning, alleviating the intricate design process of handling domain-specific messages between them. Through a series of simulations in a box-packing domain, we quantitatively demonstrate the effectiveness of LLM3 in solving TAMP problems and the efficiency in selecting action parameters. Ablation studies underscore the significant contribution of motion failure reasoning to the success of LLM3. Furthermore, we conduct qualitative experiments on a physical manipulator, demonstrating the practical applicability of our approach in real-world settings.
I Introduction
Sequential manipulation planning is an essential capability for robots to autonomously perform diverse tasks in complex environments. Executable motions must be effectively generated for robots to achieve long-term task objectives, requiring efficient planning algorithms for responsive operation and reasoning capabilities to anticipate environmental changes. Task and Motion Planning (TAMP) formulates a promising methodology that hierarchically decomposes planning into two stages: the high-level symbolic task planning stage reasons over long-horizon abstract action sequences, and the low-level continuous motion planning stage computes feasible trajectories subject to geometric constraints. In recent years, TAMP has enabled significant advances in diverse applications [1, 2, 3, 4, 5, 6]. However, a persistent challenge remains to properly interface between the task planner and the motion planner to efficiently solve TAMP, i.e., generating action sequences that satisfy both symbolic task goals and continuous motion constraints.
Figure 1: The proposed LLM3 framework. (a) Traditional TAMP frameworks rely on manually designed, domain-specific modules for interfacing between task and motion planners. (b) In contrast, we leverage a pre-trained LLM to iteratively propose refined plans and action parameters, by reasoning on motion planning failures.
Traditional TAMP approaches often rely on manually designed modules to interface between symbolic and continuous domains, as depicted in Fig. 1(a). These modules serve two key roles. First, they act as action parameter samplers that generate real-valued parameters for symbolic actions. These parameters provide numerical goals to the motion planner. Selecting appropriate action parameters, e.g., in the object rearrangement task, a suitable 2D target location $(x,y)$ for action Place(object), is crucial for the motion feasibility of actions and the efficiency of TAMP. While previous works propose to learn heuristic parameter samplers from data [7, 8], they are tailored to specific domains and lack generalizability. Second, these modules implement mechanisms to incorporate motion failure into the task planner to generate improved action plans, e.g., by updating the symbolic state [9]. However, they usually require domain-specific design by human experts. In summary, these modules are domain-specific and require substantial manual effort to design, which hinders generalizability to novel environments.
Recent Large Language Models (LLMs[10] pre-trained on web-scale text data have demonstrated emergent capabilities in reasoning [11] and planning [12] when provided with in-context prompts [13, 14]. Pre-trained LLMs can perform task planning [12, 15, 16], generate continuous parameters [17], and reason on environment feedback [18, 16]. Our intuition is that pre-trained LLMs could provide a general and domain-independent approach to interfacing between symbolic and continuous domains for TAMP, eliminating the need to design domain-specific modules manually.
In this paper, we present LLM3 (Large Language Model-based Task and Motion Planning with Motion Failure Reasoning), an LLM-powered TAMP framework that reasons over motion planning feedback for effective planningFig. 1(b). Specifically, LLM3 employs a pre-trained LLM to (i) propose symbolic action sequences towards the task goal, (ii) generate continuous action parameters that lead to feasible motion, and (iii) reason over motion planning feedback to iteratively refine the proposed symbolic actions and parameters. This framework offers several key advantages over traditional TAMP approaches. First, it does not require manually designed symbolic domain files for task planning, instead leveraging the knowledge encoded in the LLM to propose symbolic actions. Second, it uses the LLM as a domain-independent informed parameter sampler to generate continuous action parameters, which benefits from the implicit heuristics of the LLM [17]. Third, its reasoning over motion planning feedback is independent of the specific choice of planner. Crucially, we categorize and organize the possible motion planning feedback to feature two major motion failure modes, i.e., collision and unreachability. Such motion planning feedback allows LLM3 to refine the generated action sequences and parameters in a more targeted way, and find a feasible TAMP solution with fewer planning iterations and motion planning queries.
We evaluate LLM3 in a simulated tabletop box-packing task, which poses challenges in reasoning about potential failure modes, collisions, and unreachable areas, throughout the sequential manipulation planning problem. Quantitative results demonstrate the effectiveness of LLM3, with ablation studies verifying: (i) reasoning over motion feedback significantly improves success rates and planning efficiency, and (ii) the LLM-based parameter sampler is substantially more sample efficient than a random sampler. Furthermore, we conduct real-robot experiments to show that LLM3 can be applied to real-world problems.
In summary, our contributions are threefold:
  1. 1.
    We introduce LLM3, the first TAMP framework that holistically employs a pre-trained LLM as a domain-independent task planner, informed action parameter sampler, and motion failure reasoner.
  2. 2.
    We categorize and organize the feedback from the motion planner, which enables LLM3 to efficiently identify and resolve planner-independent motion failures through targeted refinement of actions and parameters.
  3. 3.
    We conduct comprehensive experiments in both the simulation and the real world to demonstrate the effectiveness of LLM3 in solving TAMP problems.
I-A Related Work
Task and Motion Planning
Traditional TAMP approaches employ a high-level task planner to generate symbolic action sequences and a low-level motion planner to generate motion trajectories. The task planner requires pre-designed symbolic planning domains represented in formatted representations, such as Planning Domain Definition Language (PDDL). Significant efforts have been made to develop manually engineered modules that interface the task planner and motion planner, such as incorporating motion-level constraints into task planning [19, 4], making approximations at the motion level [20, 21], and designing specialized communication modules [9]. However, manually defining task planning domains and interface modules to fully capture real-world complexity is impractical. Furthermore, as the action space grows, searching for geometrically feasible symbolic action sequences becomes computationally challenging without effective heuristics [19]. Recent work has explored data-driven heuristics to improve TAMP efficiency [7, 22, 23, 24], but these domain-specific heuristics lack generalizability across domains. In this work, we employ a pre-trained LLM as both the task planner and the interface between task and motion. We expect that semantic knowledge in the LLM can provide domain-independent heuristics for TAMP.
Robot Planning with LLMs
Recent Large Language Models (LLMs[10] encode vast world knowledge and exibit the emergent capability for planning [12, 25] through few-shot or zero-shot in-context learning [13, 14]. Pre-trained LLMs have been applied for task planning of robots or embodied agents [15, 16, 26, 27, 28, 29, 18, 30, 31]. Notably, Inner Monologue [16] takes in textualized environment feedback and generate actions to execute, while ReAct [18] further advanced this closed-loop approach by integrating reasoning and acting. Voyager [29] and DEPS [28] focus on developing open-ended embodied agents that iteratively replan based on execution failure in PC games. Furthermore, LLMs have been applied to solve TAMP problems  [32, 33]. Specifically, LLM-GROP [32] employs a pre-trained LLM to instantiate symbolic goals and continuous object placements for semantic object rearrangement, which are then fed into a classical TAMP planner. AutoTAMP [33] leverages an LLM to translate natural-language task specifications into a formal language processable by off-the-shelf TAMP algorithms. Our usage of LLMs in TAMP is inspired by many of the above works; however, the major difference is that we leverage the LLM as the core component of our TAMP framework.
I-B Overview
The remainder of this paper is organized as follows. We first provide preliminaries of TAMP and LLM-based planning in Section II. Then we elaborate on the proposed LLM3 framework in Section III. We summarize the results of simulated and real robot experiments and provide a discussion in Section IV. Finally, Section V concludes the paper and acknowledges the limitations of LLM3.
Figure 2: The system diagram of the proposed LLM3 framework. (a) We show an example of utilizing a pre-trained LLM for reasoning and generating action sequences. (b) The feasibility of the proposed action sequence is verified by rollout with a motion planner and transition function $\mathcal{T}$. The motion planning feedback is saved into a trace that is provided to the LLM in the next iteration.
II Preliminaries and Problem Setting
II-A Task and Motion Planning
A Task and Motion Planning (TAMP) problem [3] is a tuple $<\mathcal{O},\mathcal{S},\mathcal{A},\mathcal{T},s_{0},g>$, where:
  • $\mathcal{O}$ is the set of objects in the environment.
  • $\mathcal{S}$ is the state space factorized with respect to objects $\mathcal{O}$, where a state $s\in\mathcal{S}$ comprises the state of all objects, e.g., their 3D positions and dimensions.
  • $\mathcal{A}$ is the set of primitive actions with low-level execution handled by a motion planner. A primitive action $a\in A$ is parameterized by object variables $\bar{o}$ and continuous parameters $\theta$, which can be instantiated with specific objects $\underline{o}=(o_{1},...,o_{m})$ and parameter values to produce a ground action $\underline{a}$, e.g., $\texttt{Place}(\texttt{block},[0.1,0.2])$. The parameters $\theta$ provide the goal for the motion planner; we say $\underline{a}$ is feasible when the motion planner has a solution, i.e., a collision-free trajectory $\tau$. We denote a feasible action by $\underline{a}(\tau)$.
  • $\mathcal{T}$ is the state transition function that outputs the next state $s_{t+1}$ after executing an action $\underline{a}_{t}(\tau_{t})$ at a state $s_{t}$, i.e., $s_{t+1}=\mathcal{T}(s_{t},\underline{a}_{t}(\tau_{t}))$. We assume that $\mathcal{T}(s_{t},\underline{a}_{t}(\tau_{t})))$ can be evaluated with a black-box simulator.
  • $s_{0}$ is the initial state that follows $s_{0}\in\mathcal{S}$.
  • $g$ is the goal function $g:~{}\mathcal{S}\rightarrow\{0,1\}$, which checks whether the task goal is achieved at state $s$.
The objective of TAMP is to derive a sequence of feasible actions $(\underline{a}_{0}(\tau_{0}),\underline{a}_{1}(\tau_{2}),...,\underline{a}_{T}(\tau_{T}))$, such that $g(s_{T+1})=1$ and $s_{t+1}=\mathcal{T}(s_{t},\underline{a}_{t}(\tau_{t})))$, where $t=0,1,...,T$.
A common strategy to solve the TAMP problem is known as “search-then-sample” [9, 1, 3], which alternates generating symbolic action sequences through backtracking search and sampling the continuous action parameters until a feasible plan is found to reach the goal. However, a naive search-then-sample task and motion planner is usually inefficient, as verifying the action feasibility requires invoking the computationally expensive motion planning process. To mitigate this complexity, researchers have crafted symbolic domains [9, 34] to prune infeasible symbolic action sequences, or learned heuristic samplers to sample action parameters [8, 35] that lead to feasible plans. In this work, we leverage pre-trained LLMs as both a task planner and a heuristic sampler to generate symbolic action sequences and parameters that tend to produce feasible plans.
II-B Planning as Sequence Prediction
Recently, planning problems have been formulated as sequence prediction to avoid the computationally expensive search process [36, 37, 38]. In particular, pre-trained LLMs are employed to generate discrete actions [12, 25] and continuous parameters [26] in an auto-regressive manner when provided in-context prompts [14]. Formally, with textualized initial state $s_{0}$, goal $g$ and additional context $c$, we have:
$\underline{a}_{0:T}$ $=\arg\max~{}p_{LM}(\underline{a}_{0:T}|s_{0},g,c)$ (1)
$=\arg\max~{}p_{LM}(\underline{a}_{0}|,s_{0},g,c)\prod_{t=1}^{T}p_{LM}(\underline{a}_{t}|\underline{a}_{0:t-1},s_{0},g,c),$
where $p_{LM}$ is the generative probability of a pre-trained language model. Following this scheme, we use pre-trained LLMs to propose sequences of symbolic actions and continuous action parameters based on the initial state, goal, and trace of motion planning feedback. As the proposed action sequences may be infeasible at the motion level, we validate the feasibility of proposed action sequences with a motion planner, the state transition function $\mathcal{T}$ and the goal $g$. We iterate these two steps until a feasible plan is found.
III Method
We introduce LLM3, a TAMP framework that leverages a pre-trained LLM to reason on motion failure and generate iteratively refined symbolic actions and continuous action parameters. The system diagram of LLM3 is shown in Fig. 2. Below, we elaborate on the overall LLM3 framework, reasoning and planning with pre-trained LLM, and the designed motion planning feedback.
Input : pre-trained LLM with prompt template to generate action sequences $LLM$, state transition function $\mathcal{T}$ with motion planner $MP$, goal function $g$, initial state $s_{0}$, maximum number of planning attempts $N_{max}$, maximum trace size $k$
Output : success indicator $success$, a sequence of feasible actions $plan$
// initialize output and feedback trace $plan\leftarrow[]$, $success\leftarrow\text{False}$, $feedback\leftarrow[]$, $trace\leftarrow[]$ // initialize current state and planning iteration count $s\leftarrow s_{0}$, $iter\leftarrow 0$ // main planning loop while not $success$ and $iter<N_{max}$ do
       // reason and plan with LLM $reason,llm\_plan\leftarrow LLM(s_{0},trace)$, $iter\leftarrow iter+1$ // rollout $llm\_plan$ from $s$ with motion planner foreach $\underline{a}\in llm\_plan$ do
             // call motion planner to verify feasibility of a $\tau,mp\_feedback\leftarrow MP(s;\underline{a})$ $feedback.append((\underline{a},mp\_feedback))$ // terminate rollout if action $\underline{a}$ is infeasible if $\tau$ is $None$ then
1                   break
            // update state $s$ and add $\underline{a}$ to $plan$ $s\leftarrow\mathcal{T}(s,\underline{a}(\tau))$, $plan.append(\underline{a}(\tau))$
      // check whether $s$ satisfies goal $success,task\_feedback\leftarrow g(s)$ // when the plan is feasible but doesn’t reach goal if not $success$ then
             // record $task\_feedback$ $feedback.append(task\_feedback)$ // add aggregated feedback to $trace$ $trace.append(feedback)$, $trace\leftarrow trace[-k:]$ // clear aggregated feedback $feedback\leftarrow[]$
2
return $success$, $plan$
Algorithm 1 LLM3 for TAMP
III-A The LLM3 Framework
As shown in Algorithm Algorithm 1 and Figure 2, the LLM3 framework iterates between: (i) reasoning on previous motion failure and generating an action sequence (i.e., symbolic actions and continuous parameters) with a pre-trained LLM, and (ii) verifying the feasibility of the action sequence with a motion planner. At each planning iteration, the LLM takes the current state $s$ and the trace of motion planning feedback $trace$, and outputs the reasoning for the previous motion failure $reason$ and an action sequence $llm\_plan$ to solve the TAMP problem (see Section III-B). The motion planner then attempts to find a collision-free motion trajectory for each action $\underline{a}\in llm\_plan$ sequentially. We synthesize motion planning feedback (see Section III-C) for each motion planner query and aggregated the feedback for all actions in $llm\_plan$. The planning iteration ends with failure when an action has no feasible motion or the action sequence fails to reach the goal. The aggregated feedback is then added to a trace $trace$, which is maintained throughout the life cycle of the framework with a maximum size $k$. In the next planning iteration, the trace is fed into the LLM to generate motion failure reasoning and another action sequence that improves on the previous one. This process repeats until a generated action sequence reaches the goal with no motion failure or the maximum number of attempts is exceeded. Overall, the LLM3 framework can be regarded as a search-then-sample TAMP planner that generates action sequences with incrementally improved quality, guided by the intrinsic heuristics of the pre-trained LLM and the previous motion failure. By design, we expect LLM3 to exhibit superior efficiency compared to unguided planners that sample action parameters randomly.
III-B Reasoning and Planning with pre-trained LLM
Following previous attempts in utilizing LLMs for reasoning and planning [11, 12, 16, 18], we prompt a pre-trained LLM to generate motion failure reasoning and action sequences in text format. Since we want to limit the domain-specific prior provided to the LLM, we use zero-shot prompting [11] without providing any planning examples. We adopt Chain-of-Thought (CoT) prompting [39, 18] to have the LLM generate the reasoning and action sequence in an auto-regressive manner. The prompt fed into LLM comprises the following contents: (i) a system message that provides the global context to the pre-trained LLM and activates its planning capability, (ii) a task description specifying the environment, goal, and available primitive actions, (iii) the textualized initial environment state, (iv) the trace of motion planning feedback, and (v) the output format. Note that most promp content will remain unchanged for different planning iterations of the same TAMP problem. Only the motion planning feedback trace in (iv) will be updated as new failed action sequences are added. Fig. 2(a) illustrated an example of reasoning and planning with LLM.
We implement two strategies for the pre-trained LLM to generate a new action sequence that improves on the previous one: (i) backtrack, where we expect the LLM to backtrack to a previous action that has feasible motion, and continual to generate actions that complete the plan, and (ii) from scratch, where we expect the LLM to directly generate a new action sequence that attempts to avoid the motion failure happened to its previous output. We achieve this by designing the system message and output format description in the prompt. We show the prompt template for the two variants in Fig. 3, where the content specific to each variant is highlighted.
[…]
Figure 3: Prompt templetes used by LLM3. We show alternative contents specific for the backtrack variant in orange and from scratch variant in blue.
III-C Synthesizing Motion Planning Feedback
We realize a ground action $\underline{a}$ by calculating a collision-free trajectory $\tau$ with a sampling-based motion planner, e.g. Bi-directional Rapidly-exploring Random Trees (BiRRT) [40]. The input to the motion planner includes the initial environment state (includes the robot state), and a goal pose of the robot end effector specified by the continuous action parameters of $\underline{a}$. The motion planner samples and searches for a collision-free joint space trajectory for the robot to reach the goal end-effector pose.
By default, the motion planner reports a binary signal that indicates whether there is a feasible trajectory. It does not give more abstract-level feedback, which explains why motion planning fails. As a result, the TAMP planners can acquire useful feedback from motion planning failures to improve high-level planning. To this end, we additionally synthesize semantically meaningful motion-level feedback so that LLM3 can improve on previous failures more effectively. We observe that typical motion planning failures can be categorized into two types, i.e., collisions and unreachability. Therefore, we synthesize categorized motion planning feedback following the templates below (Fig. 4):
Figure 4: Three types of motion possibilities. A: the object placement is in collision with an existing object. B: the object placement is beyond the robot’s reach. C: the object placement is feasible.
  1. (A)
    The goal configuration is in collision with object.
  2. (B)
    The goal configuration has no feasible inverse kinematics (IK) solution.
  3. (C)
    The goal configuration is collision-free and reachable.
In practice, we integrate the motion planner with an additional IK solver and collision checker for obtaining these feedbacks, finding this design to be practically effective.
IV Simulation and Experiment
In simulations, we initially perform an ablation study on our LLM3 framework in two settings of the tabletop box-packing task, quantitatively evaluating its effectiveness based on i) the planning success rate (%SR), ii) the number of LLM calls (#LM), iii) and the number of motion planner calls (#MP). Additionally, we demonstrate the role of LLM as an informed action parameter sampler by comparing it to a baseline utilizing random sampling strategies. Finally, we validate the proposed LLM3 framework through experimentation on a perception-integrated physical robotic manipulator, confirming its validity in real-world scenarios.
IV-A Simulation Setup
We developed a PyBullet-based simulation environment for our box-packing tasks, as illustrated in Fig. 5a. In Setting 1, three different sets of objects are given with increasing total sizes, while the size of the basket remains constant. This task requires the LLM3 to oversee potential collisions between objects or the robot throughout the action sequence. The LLM must reason why collisions occur and adjust previous actions to ensure feasible task and motion plans. Setting 2 involves placing the Set 3 objects into baskets of increasing sizes. Here, the robot cannot access the entire basket region but encounters a collision likelihood similar to the most crowded condition in Setting 1. In this setting, the task becomes more complex as it challenges the LLM to reason about the robot’s operational space and adjust the plan accordingly. Throughout the simulations, we utilize GPT-4 Turbo [41] as the LLM planner and BiRRT [40] as the motion planner, with 10 attempts for each setting. The code for replicating the experiment is available at https://github.com/AssassinWS/LLM-TAMP.
Figure 5: The box-packing task setup in a simulated environment. (a) The task requires the robot to place one of (b) three sets of objects fully into the basket. (c) In setting 1, the total object size increases but the basket sizes remain the same. All baskets are fully reachable by the robot. (d) In setting 2, the basket size increases, but some portions of baskets are longer within the robot’s reach.
IV-B Ablation Study
The conducted ablation study compares the proposed LLM3 with baseline methods:
TABLE I: Ablation Study
Method Setting 1 Setting 2
Easy Medium Hard Small Medium Large
%SR #LM #MP %SR #LM #MP %SR #LM #MP %SR #LM #MP %SR #LM #MP %SR #LM #MP
LLM3 Backtrack 100 1.6 11.8 100 4.4 28.4 60 11.4 39.8 60 11.4 39.8 80 9.5 50 50 13.5 44.8
Backtrack 100 1.8 12.6 90 6.3 32 40 15.1 55.3 40 15.1 55.3 30 14.6 16 30 15.8 48
LLM3 Scratch 100 1.7 13.2 100 7 46.1 70 8.8 30.9 70 8.8 30.9 70 11.5 50.2 60 10.6 32
Scratch 100 2.4 17.6 60 12.3 45.3 50 13.7 42.8 50 13.7 42.8 30 16.2 45.7 40 13.2 24
  1. 1.
    LLM3 Backtrack: The proposed LLM3 framework backtrack variant. See Fig. 3
  2. 2.
    Backtrack: The LLM proposes plans with backtracking but without motion planning feedback (line 7).
  3. 3.
    LLM3 Scratch: The proposed LLM3 framework from scratch variant. It replans the entire action sequence if any action fails, incorporating motion planning feedback. See Fig. 3.
  4. 4.
    Scratch: The LLM plans the action sequence once and executes the plan without any feedback.
Three evaluation criteria are considered: The number of LLM calls (#LM) counts how many times the LLM API is called during planning. A lower #LM indicates that the planner can produce a feasible task plan more efficiently. In each attempt, #LM has a maximum cap of 20; attempts with over 20 #LMs will be counted as a failure. The total success rate is recorded as %SR. Additionally, the number of motion planner calls, #MP, is another critical criterion as in traditional TAMP approaches, where massive and time-consuming motion planner calls are the main cause for their inefficiency. The study results are summarized in Table I.
In both settings, integrating motion planning feedback results in a decrease in #LM and #MP, along with an increase in %SR for both backtrack and scratch strategies. This indicates that the LLM can reason about failures from motion planning feedback, and importantly, propose adjusted task plans and action parameters that are more likely to produce feasible motions. Surprisingly, there is no clear evidence to suggest that utilizing backtracking is superior to replanning from scratch. We’ve observed distinct behaviors among different strategies employed by LLM3. Specifically, when utilizing backtracking, LLM sometimes consistently adjusts the action parameter of the specific action that directly led to failure, without adjusting actions several steps prior. In contrast, the LLM3 employing a planning from scratch strategy simply re-samples all actions and parameters, occasionally resulting in a slightly lower #LM.
IV-C Action Parameter Selection
TABLE II: Comparison of Different Sampling Strategies
Method #Itration #MP
Random Sampling 109.6 663.1
LLM Sampling 10.8 70.2
LLM Sampling + Feedback 7.9 53.2
This study examines whether an LLM can function as an informed action parameter sampler, using the Medium setup from Setting 1 in the previous section. In this study, the action sequence is predetermined and consists of a total of 8 steps, where the robot sequentially performs pick and place actions to relocate all four blocks into the basket. However, the specific placement location is not provided, requiring the action parameters to be sampled to ensure the feasibility at the motion level. We implement three methods, each runs 50 attempts, for comparison:
  1. 1.
    Random Samples: A random sampler is implemented to independently and uniformly sample all the block placement locations within the basket region.
  2. 2.
    LLM: The LLM is provided with the task setting and the symbolic action sequence, prompting it to select the action parameters to make the action sequence feasible.
  3. 3.
    LLM + Feedback: Building upon 2), the LLM is additionally provided with motion planning feedback if the previous attempt fails.
The results summarized in Table II demonstrate the efficacy of using LLM as an informed action parameter sampler in the context of box-packing tasks. Specifically, while random sampling requires an average of 109.6 iterations and 663.1 #MP to achieve feasible action sequences, the LLM substantially reduces them to an average of 10.8 iterations and 70.2 #MP. When incorporating motion planning feedback alongside the LLM, the sampling requirement decreases even further to an average of 7.9 iterations and 53.2 #MP. These findings underscore the ability of LLMs to efficiently select action parameters that align with task constraints, resulting in a notable reduction in the number of motion planning calls needed to generate feasible action sequences. Moreover, the additional benefit gained from integrating motion planning feedback highlights the importance of incorporating real-time feedback mechanisms to further refine and optimize the sampling process. Overall, these results highlight the potential of LLMs as a valuable tool to improve the efficiency and effectiveness of robotic manipulation tasks, particularly in scenarios where efficient action parameter selection is crucial for improving planning performance.
IV-D Experiment Setup
Figure 6: The real-world experiment on a physical robot. The figure to the left shows the box-packing task setup. Actions 1 to 8 are proposed by LLM3 and successfully carried out by the physical manipulator.
To validate the effectiveness of our proposed method in a real-world setting, we conducted an experiment using a Franka Research 3 manipulator. The goal was to demonstrate the robot’s ability to perceive and manipulate objects in an environment with uncertainties in both perception and execution. The robot observed a single point cloud from a third-person-view RGB-D camera (Kinect Azure), capturing the workspace containing various objects such as blocks and a plate. To identify and locate individual objects, we employed Grounded Segment Anything (Grounded-SAM) [42] for object segmentation, initially segmenting objects in the 2D RGB image and then projecting the results onto the corresponding 3D point cloud. This approach yielded per-object point clouds, essential for planning and executing manipulation tasks.
Fig. 6 presents a qualitative evaluation of our method, where the robot was tasked with placing all blocks on the plate. The results demonstrate that our method enabled the robot to successfully identify and manipulate objects despite the uncertainties and challenges in the cluttered environment. The successful execution of this experiment validates the practicality and robustness of our approach, showcasing its potential for various real-world applications such as object sorting, assembly tasks, and household assistance.
V Conclusions
In this paper, we introduced LLM3, a new TAMP framework powered by the Large Language Model (LLM). Unlike conventional TAMP frameworks, which typically entail significant efforts to explicitly design domain-specific symbolic task planners and manually interface them with motion planners, LLM3 leverages the rich knowledge encoded in and the powerful reasoning capability processed by LLMs to (i) propose action sequences based on task goals without requiring a prior planning domain, (ii) automatically generate continuous action parameters for the robot motion planner, and more importantly, (iii) refine task and/or motion plans in response to motion planning failures.
Following validation through various simulations and experiments, we demonstrated that LLM3 effectively produced and refined task and motion plans for box-packing problems, exhibiting promising potential in addressing previously unspecified tasks. Our study also revealed that although the LLM can generate action parameters more efficiently than random samplers, it still necessitated multiple feedback iterations and motion planner calls. Looking ahead, the incorporation of in-context learning or fine-tuning techniques holds promise for further enhancing its efficiency, representing a crucial step towards empowering robots to tackle emerging tasks in real-world scenarios.
Acknowledgement: We thank Hang Li, Lecheng Ruan, Junyi Liang for their fruitful discussions and support. This work was supported in part by the National Natural Science Foundation of China (No.52305007).

### v3

@@
 Abstract
-Conventional Task and Motion Planning (TAMP) approaches rely on manually crafted interfaces connecting symbolic task planning with continuous motion generation.
+Conventional Task and Motion Planning (TAMP) approaches rely on manually designed interfaces connecting symbolic task planning with continuous motion generation.
 These domain-specific and labor-intensive modules are limited in addressing emerging tasks in real-world settings.
@@
 Executable motions must be effectively generated for robots to achieve long-term task objectives, requiring efficient planning algorithms for responsive operation and reasoning capabilities to anticipate environmental changes.
-Task and Motion Planning (TAMP) formulates a promising methodology that hierarchically decomposes planning into two stages: the high-level symbolic task planning stage reasons over long-horizon abstract action sequences, and the low-level continuous motion planning stage computes feasible trajectories subject to geometric constraints.
-In recent years, TAMP has enabled significant advances in diverse applications [1, 2, 3, 4, 5, 6].
+Task and Motion Planning (TAMP) formulates a methodology that hierarchically decomposes planning into two stages: the high-level symbolic task planning stage reasons over long-horizon abstract action sequences, and the low-level continuous motion planning stage computes feasible trajectories subject to geometric constraints.
+In recent years, TAMP has enabled significant advances [1, 2, 3, 4, 5, 6, 7].
 However, a persistent challenge remains to properly interface between the task planner and the motion planner to efficiently solve TAMP, i.e., generating action sequences that satisfy both symbolic task goals and continuous motion constraints.
@@
 In summary, these modules are domain-specific and require substantial manual effort to design, which hinders generalizability to novel environments.
-Recent Large Language Models (LLMs[10] pre-trained on web-scale text data have demonstrated emergent capabilities in reasoning [11] and planning [12] when provided with in-context prompts [13, 14].
-Pre-trained LLMs can perform task planning [12, 15, 16], generate continuous parameters [17], and reason on environment feedback [18, 16].
+Recent Large Language Models (LLMs) pre-trained on web-scale text data have demonstrated emergent capabilities in reasoning [11] and planning [12].
+Pre-trained LLMs can perform task planning [13], generate continuous parameters [14], and reason on environment feedback [13].
 Our intuition is that pre-trained LLMs could provide a general and domain-independent approach to interfacing between symbolic and continuous domains for TAMP, eliminating the need to design domain-specific modules manually.
@@
 Task and Motion Planning
-Traditional TAMP approaches employ a high-level task planner to generate symbolic action sequences and a low-level motion planner to generate motion trajectories.
+Conventional TAMP approaches employ a high-level task planner to generate symbolic action sequences and a low-level motion planner to generate motion trajectories.
 The task planner requires pre-designed symbolic planning domains represented in formatted representations, such as Planning Domain Definition Language (PDDL).
@@
 Furthermore, as the action space grows, searching for geometrically feasible symbolic action sequences becomes computationally challenging without effective heuristics [15].
-Recent work has explored data-driven heuristics to improve TAMP efficiency [7, 22, 23, 24], but these domain-specific heuristics lack generalizability across domains.
+Recent work has explored data-driven heuristics to improve TAMP efficiency [8, 18], but these domain-specific heuristics lack generalizability across domains.
 In this work, we employ a pre-trained LLM as both the task planner and the interface between task and motion.
@@
 Robot Planning with LLMs
-Recent Large Language Models (LLMs[10] encode vast world knowledge and exibit the emergent capability for planning [12, 25] through few-shot or zero-shot in-context learning [13, 14].
-Pre-trained LLMs have been applied for task planning of robots or embodied agents [15, 16, 26, 27, 28, 29, 18, 30, 31].
+Recent Large Language Models (LLMs[19, 20] encode vast world knowledge and exibit the emergent capability for planning [12].
+Pre-trained LLMs have been applied for task planning of robots or embodied agents [13, 21, 22, 23, 24, 25, 26, 27].
 Notably, Inner Monologue [13] takes in textualized environment feedback and generate actions to execute, while ReAct [28] further advanced this closed-loop approach by integrating reasoning and acting.
-Voyager [29] and DEPS [28] focus on developing open-ended embodied agents that iteratively replan based on execution failure in PC games.
+Voyager [22] and DEPS [21] focus on developing open-ended embodied agents that iteratively replan based on execution failure.
 Furthermore, LLMs have been applied to solve TAMP problems  [29, 30].
@@
 Our usage of LLMs in TAMP is inspired by many of the above works; however, the major difference is that we leverage the LLM as the core component of our TAMP framework.
-I-B Overview
-The remainder of this paper is organized as follows.
-We first provide preliminaries of TAMP and LLM-based planning in Section II.
-Then we elaborate on the proposed LLM3 framework in Section III.
-We summarize the results of simulated and real robot experiments and provide a discussion in Section IV.
-Finally, Section V concludes the paper and acknowledges the limitations of LLM3.
-Figure 2: The system diagram of the proposed LLM3 framework.
+Figure 2: System diagram of the proposed LLM3 framework.
 (a) We show an example of utilizing a pre-trained LLM for reasoning and generating action sequences.
@@
 II-A Task and Motion Planning
-A Task and Motion Planning (TAMP) problem [3] is a tuple $<\mathcal{O},\mathcal{S},\mathcal{A},\mathcal{T},s_{0},g>$, where:
+A TAMP problem is a tuple $<\mathcal{O},\mathcal{S},\mathcal{A},\mathcal{T},s_{0},g>$, where:
   • $\mathcal{O}$ is the set of objects in the environment.
@@
 A common strategy to solve the TAMP problem is known as “search-then-sample” [10, 1, 3], which alternates generating symbolic action sequences through backtracking search and sampling the continuous action parameters until a feasible plan is found to reach the goal.
-However, a naive search-then-sample task and motion planner is usually inefficient, as verifying the action feasibility requires invoking the computationally expensive motion planning process.
-To mitigate this complexity, researchers have crafted symbolic domains [9, 34] to prune infeasible symbolic action sequences, or learned heuristic samplers to sample action parameters [8, 35] that lead to feasible plans.
-In this work, we leverage pre-trained LLMs as both a task planner and a heuristic sampler to generate symbolic action sequences and parameters that tend to produce feasible plans.
+However, a naive search-then-sample task and motion planner is usually inefficient, as verifying the action feasibility requires invoking the computationally expensive motion planning process [31, 32].
+To mitigate this complexity, researchers have crafted symbolic domains [10, 33, 34] to prune infeasible symbolic action sequences, or learned heuristic samplers to sample action parameters [9, 35] that lead to feasible plans.
+In this work, we leverage pre-trained LLMs as both a task planner and a heuristic sampler to generate symbolic action sequences and action parameters.
 II-B Planning as Sequence Prediction
-Recently, planning problems have been formulated as sequence prediction to avoid the computationally expensive search process [36, 37, 38].
-In particular, pre-trained LLMs are employed to generate discrete actions [12, 25] and continuous parameters [26] in an auto-regressive manner when provided in-context prompts [14].
+Recently, planning problems have been formulated as sequence prediction to avoid the computationally expensive search process [36].
+In particular, pre-trained LLMs are employed to generate discrete actions [12] and continuous parameters in an auto-regressive manner when provided in-context prompts.
 Formally, with textualized initial state $s_{0}$, goal $g$ and additional context $c$, we have:
@@
 2.
-Below, we elaborate on the overall LLM3 framework, reasoning and planning with pre-trained LLM, and the designed motion planning feedback.
+Below, we elaborate on the overall framework, reasoning and planning with the pre-trained LLM, and the designed motion planning feedback.
+1
 Input : pre-trained LLM with prompt template to generate action sequences $LLM$, state transition function $\mathcal{T}$ with motion planner $MP$, goal function $g$, initial state $s_{0}$, maximum number of planning attempts $N_{max}$, maximum trace size $k$
 Output : success indicator $success$, a sequence of feasible actions $plan$
-// initialize output and feedback trace $plan\leftarrow[]$, $success\leftarrow\text{False}$, $feedback\leftarrow[]$, $trace\leftarrow[]$ // initialize current state and planning iteration count $s\leftarrow s_{0}$, $iter\leftarrow 0$ // main planning loop while not $success$ and $iter<N_{max}$ do
-       // reason and plan with LLM $reason,llm\_plan\leftarrow LLM(s_{0},trace)$, $iter\leftarrow iter+1$ // rollout $llm\_plan$ from $s$ with motion planner foreach $\underline{a}\in llm\_plan$ do
-             // call motion planner to verify feasibility of a $\tau,mp\_feedback\leftarrow MP(s;\underline{a})$ $feedback.append((\underline{a},mp\_feedback))$ // terminate rollout if action $\underline{a}$ is infeasible if $\tau$ is $None$ then
+2
+// initialize output and feedback trace
+3 $plan\leftarrow[]$, $success\leftarrow\text{False}$, $feedback\leftarrow[]$, $trace\leftarrow[]$
+// initialize current state and planning iteration count
+4 $s\leftarrow s_{0}$, $iter\leftarrow 0$
+// main planning loop
+5 while not $success$ and $iter<N_{max}$ do
+       // reason and plan with LLM
+6       $reason,llm\_plan\leftarrow LLM(s_{0},trace)$, $iter\leftarrow iter+1$
+       // rollout $llm\_plan$ from $s$ with motion planner
+7       foreach $\underline{a}\in llm\_plan$ do
+             // call motion planner to verify feasibility of a
+8             $\tau,mp\_feedback\leftarrow MP(s;\underline{a})$
+9             $feedback.append((\underline{a},mp\_feedback))$
+             // terminate rollout if action $\underline{a}$ is infeasible
+10             if $\tau$ is $None$ then
 11                   break
-            // update state $s$ and add $\underline{a}$ to $plan$ $s\leftarrow\mathcal{T}(s,\underline{a}(\tau))$, $plan.append(\underline{a}(\tau))$
-      // check whether $s$ satisfies goal $success,task\_feedback\leftarrow g(s)$ // when the plan is feasible but doesn’t reach goal if not $success$ then
-             // record $task\_feedback$ $feedback.append(task\_feedback)$ // add aggregated feedback to $trace$ $trace.append(feedback)$, $trace\leftarrow trace[-k:]$ // clear aggregated feedback $feedback\leftarrow[]$
-2
-return $success$, $plan$
+12
+            // update state $s$ and add $\underline{a}$ to $plan$
+13             $s\leftarrow\mathcal{T}(s,\underline{a}(\tau))$, $plan.append(\underline{a}(\tau))$
+14
+      // check whether $s$ satisfies goal
+15       $success,task\_feedback\leftarrow g(s)$
+       // when the plan is feasible but doesn’t reach goal
+16       if not $success$ then
+             // record $task\_feedback$
+17             $feedback.append(task\_feedback)$
+             // add aggregated feedback to $trace$
+18             $trace.append(feedback)$, $trace\leftarrow trace[-k:]$
+             // clear aggregated feedback
+19             $feedback\leftarrow[]$
+20
+21
+22return $success$, $plan$
 Algorithm 1 LLM3 for TAMP
@@
 III-B Reasoning and Planning with pre-trained LLM
-Following previous attempts in utilizing LLMs for reasoning and planning [11, 12, 16, 18], we prompt a pre-trained LLM to generate motion failure reasoning and action sequences in text format.
-Since we want to limit the domain-specific prior provided to the LLM, we use zero-shot prompting [11] without providing any planning examples.
-We adopt Chain-of-Thought (CoT) prompting [39, 18] to have the LLM generate the reasoning and action sequence in an auto-regressive manner.
+Following previous attempts in utilizing LLMs for reasoning and planning [11, 12, 13], we prompt a pre-trained LLM to generate motion failure reasoning and action sequences in text format.
+Since we want to limit the domain-specific prior provided to the LLM, we use zero-shot prompting without providing any planning examples.
+We adopt Chain-of-Thought (CoT) prompting to have the LLM generate the reasoning and action sequence in an auto-regressive manner.
 The prompt fed into LLM comprises the following contents: (i) a system message that provides the global context to the pre-trained LLM and activates its planning capability, (ii) a task description specifying the environment, goal, and available primitive actions, (iii) the textualized initial environment state, (iv) the trace of motion planning feedback, and (v) the output format.
@@
 3, where the content specific to each variant is highlighted.
-[…]
+System message: You are an AI robot that generates a plan of actions to reach the goal… You are expected to correct the plan incrementally (on top of the last plan) to avoid motion failure.
+This may involve sample new parameters for the failed action or reverse one or more succeeded actions for backtracking… You are expected to generate a plan from scratch.
+Task description: A robot arm is tasked to pack boxes into a basket on a table.
+The robot sits at (0, 0), and faces the positive x-axis, while the positive z-axis points up.
+The robot is equipped with primitive actions, each taking a list of objects and continuous parameters as input:
+$\texttt{pick}([\texttt{obj}],\{\})$: pick up obj, with no parameters.
+$\texttt{place}([\texttt{obj}],\{``x``:[0.0,1.0],``y``:[-1.0,1.0],\\ ``theta``:[-3.14,3.14]\})$: place obj at location $(x,y)$ with the planar rotation theta, where $x$ ranges $(0.0,1.0)$, $y$ ranges $(-1.0,1.0)$, and $theta$ ranges $(-3.14,3.14)$.
+Initial state: {init_state}
+motion planning feedback trace: {trace}
+Output format: Please generate output step-by-step, which includes your reasoning for the failure of the last plan as well as the generated plan… Answer the questions: (i) what is the cause of the failure of the last plan?
+(ii) can altering action parameters for the failed action solve the problem… (iii) do we need to reverse one or more succeeded actions executed before the failed action… (ii) what is your strategy to generate a new plan from scratch to accomplish the task goal?
+Please organize the output following the JSON format below: { “Reasoning”: “My reasoning for the failure of the last plan is …”, “Full Plan”: [“pick([’red_box’], {})”, “place([’red_box’], {’x’: 0.51, ’y’: 0.02, ’theta’: 0.00})”, …] }
 Figure 3: Prompt templetes used by LLM3. We show alternative contents specific for the backtrack variant in orange and from scratch variant in blue.
@@
   2. (B)
-    The goal configuration has no feasible inverse kinematics (IK) solution.
+    The goal configuration has no feasible IK solution.
   3. (C)
@@
 IV-A Simulation Setup
+The simulation has an important role in Robotics[38, 39].
 We developed a PyBullet-based simulation environment for our box-packing tasks, as illustrated in Fig.
 5a.
-In Setting 1, three different sets of objects are given with increasing total sizes, while the size of the basket remains constant.
-This task requires the LLM3 to oversee potential collisions between objects or the robot throughout the action sequence.
+In Setting 1, three different sets of objects are given with increasing total sizes, while the basket size remains unchanged.
+This task requires the LLM3 to oversee potential collisions among objects and the robot throughout the action sequence.
 The LLM must reason why collisions occur and adjust previous actions to ensure feasible task and motion plans.
@@
 In this setting, the task becomes more complex as it challenges the LLM to reason about the robot’s operational space and adjust the plan accordingly.
-Throughout the simulations, we utilize GPT-4 Turbo [41] as the LLM planner and BiRRT [40] as the motion planner, with 10 attempts for each setting.
-The code for replicating the experiment is available at https://github.com/AssassinWS/LLM-TAMP.
+Throughout the simulations, we utilize GPT-4 Turbo as the LLM planner and BiRRT [37] as the motion planner, with 10 attempts for each setting.
+[^1] The code is available at https://github.com/AssassinWS/LLM-TAMP.
 Figure 5: The box-packing task setup in a simulated environment.
@@
   3. 3.
-    LLM3 Scratch: The proposed LLM3 framework from scratch variant. It replans the entire action sequence if any action fails, incorporating motion planning feedback. See Fig. 3.
+    LLM3 Scratch: The proposed LLM3 framework from scratch variant. It replans the entire action sequence, incorporating feedback from the motion planner if any action fails. See Fig. 3
   4. 4.
@@
 This indicates that the LLM can reason about failures from motion planning feedback, and importantly, propose adjusted task plans and action parameters that are more likely to produce feasible motions.
-Surprisingly, there is no clear evidence to suggest that utilizing backtracking is superior to replanning from scratch.
+Surprisingly, no clear evidence suggests that utilizing backtracking is superior to replanning from scratch.
 We’ve observed distinct behaviors among different strategies employed by LLM3.
-Specifically, when utilizing backtracking, LLM sometimes consistently adjusts the action parameter of the specific action that directly led to failure, without adjusting actions several steps prior.
+Specifically, when utilizing backtracking, LLM sometimes consistently adjusts the action parameter of the specific action that directly led to failure, without adjusting previous actions.
 In contrast, the LLM3 employing a planning from scratch strategy simply re-samples all actions and parameters, occasionally resulting in a slightly lower #LM.
@@
 Method #Itration #MP
-Random Sampling 109.6 663.1
-LLM Sampling 10.8 70.2
-LLM Sampling + Feedback 7.9 53.2
+Random Samples 109.6 663.1
+LLM 10.8 70.2
+LLM + Feedback 7.9 53.2
 This study examines whether an LLM can function as an informed action parameter sampler, using the Medium setup from Setting 1 in the previous section.
@@
 The goal was to demonstrate the robot’s ability to perceive and manipulate objects in an environment with uncertainties in both perception and execution.
-The robot observed a single point cloud from a third-person-view RGB-D camera (Kinect Azure), capturing the workspace containing various objects such as blocks and a plate.
-To identify and locate individual objects, we employed Grounded Segment Anything (Grounded-SAM) [42] for object segmentation, initially segmenting objects in the 2D RGB image and then projecting the results onto the corresponding 3D point cloud.
+The robot observed a single point cloud from a third-person-view RGB-D camera, capturing the workspace containing various objects such as blocks and a plate.
+To identify and locate individual objects, we employed Grounded Segment Anything [40] for object segmentation, initially segmenting objects in the 2D RGB image and then projecting the results onto the corresponding 3D point cloud.
 This approach yielded per-object point clouds, essential for planning and executing manipulation tasks.
@@
 V Conclusions
-In this paper, we introduced LLM3, a new TAMP framework powered by the Large Language Model (LLM).
-Unlike conventional TAMP frameworks, which typically entail significant efforts to explicitly design domain-specific symbolic task planners and manually interface them with motion planners, LLM3 leverages the rich knowledge encoded in and the powerful reasoning capability processed by LLMs to (i) propose action sequences based on task goals without requiring a prior planning domain, (ii) automatically generate continuous action parameters for the robot motion planner, and more importantly, (iii) refine task and/or motion plans in response to motion planning failures.
+In this paper, we introduced LLM3, a new TAMP framework powered by the pre-trained LLM.
+LLM3 leverages the rich knowledge encoded in and the powerful reasoning capability processed by LLMs to (i) propose action sequences based without requiring a prior planning domain, (ii) generate continuous action parameters for the robot motion planner, and more importantly, (iii) refine task and/or motion plans in response to motion planning failures.
 Following validation through various simulations and experiments, we demonstrated that LLM3 effectively produced and refined task and motion plans for box-packing problems, exhibiting promising potential in addressing previously unspecified tasks.
-Our study also revealed that although the LLM can generate action parameters more efficiently than random samplers, it still necessitated multiple feedback iterations and motion planner calls.
+Our study also revealed that although the pre-trained LLM can generate action parameters more efficiently than random samplers, it still necessitated multiple feedback iterations and motion planner calls.
 Looking ahead, the incorporation of in-context learning or fine-tuning techniques holds promise for further enhancing its efficiency, representing a crucial step towards empowering robots to tackle emerging tasks in real-world scenarios.
-Acknowledgement: We thank Hang Li, Lecheng Ruan, Junyi Liang for their fruitful discussions and support.
-This work was supported in part by the National Natural Science Foundation of China (No.52305007).